Mary is now four times as old as her daughter. Four years ago,the product of their ages was 40. Find their present ages.
Answer by mananth(16946)


Mary is now four times as old as her daughter. Four years ago,the product of their ages was 40. Find their present ages.
If daughter is x
Mother's age = 4x
4 years ago
daughter's age = x-4
Mother's age = 4x-4
Product of ages (x-4)((4x-4)
(x-4)((4x-4)=40
expand
4x^2-4x-16x+16 =40
4x^2-20x-24=0
x^2-5x-6=0
x^2-6x+x-6=0
x(x-6)+1(x-6)=0
(x-6)(x+1)=0
x=6 or -1
age cannot be negative
x=6
continue